OpenStudy (anonymous):

A bicycle tire has a diameter of 17 in how many revolutions of the tire are necessary to travel 800 ft

ganeshie8 (ganeshie8):

lets say "n" revolutions are needed

OpenStudy (anonymous):

800/17

ganeshie8 (ganeshie8):

nope, each revolution gives u how many feet ?

ganeshie8 (ganeshie8):

circumference = pi * d = pi * 17 inches right ?

OpenStudy (anonymous):

i got 180

ganeshie8 (ganeshie8):

to travel 800 ft = 800 * 12 inches, u wud need to make : \(\large \frac{800*12}{\pi * 17}\) revolutions

ganeshie8 (ganeshie8):

simplify

OpenStudy (anonymous):

oh dear yeah 180

ganeshie8 (ganeshie8):

180 is correct !

OpenStudy (anonymous):

thanks again :) im so bad at math

ganeshie8 (ganeshie8):

lol you're very good , u wlc :)